# Inclusive, Differentially Private Federated Learning for Clinical Data

### v2

@@ Abstract
Federated Learning (FL) offers a promising approach for training clinical AI models without centralizing sensitive patient data, yet its real-world adoption is hindered by challenges in privacy, resource constraints, and compliance. Existing differential privacy (DP) approaches often apply uniform noise, which disproportionately degrades model performance even among well-compliant institutions. In this work, we propose a novel compliance-aware FL framework that enhances DP by adaptively adjusting noise based on quantifiable client compliance scores. Additionally, we introduce a compliance scoring tool based on key healthcare and security standards to promote secure, inclusive, and equitable participation across diverse clinical settings. Extensive experiments on the public datasets demonstrate that integrating under-resourced, less compliant clinics with highly regulated institutions yields accuracy improvements of up to 15% over traditional FL. This work advances FL by balancing privacy, compliance, and performance, making it a viable solution for real-world clinical workflows in global healthcare.
Keywords:
Compliance-Aware Clinical Federated Learning Privacy-Preserving FL Adaptive ComplianceResource-Efficient DP.
@@ 2 Methods
where $n$ is the total number of compliance factors, $w_{i}$ is the weight assigned to factor $i$, and $s_{i}$ is the selected option score for factor $i$. For instance, the anonymization practices factor offers three options: ISO/TS 25237:2017 Fully Anonymized (Score 1.0), Pseudonymized (Partial Anonymization) (Score 0.7), and No Anonymization (Score 0.5), with the tool defaulting to a $0.5$ threshold, adjustable by experiment owners, including setting it to 00 if needed.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Noise Multiplier Calculation. To implement DP adaptively, noise levels are dynamically adjusted based on client compliance scores. The noise multiplier ($N_{m}$) is computed as: $N_{m}=(1.0-S_{c})+\text{Min Noise Multiplier}$, where $S_{c}$ denotes the client’s compliance score, and Min Noise Multiplier (set to $-101e\text{-}$ 10 in this experiment) ensures baseline privacy. This approach ensures that clients with lower compliance scores need higher noise levels. Noise can be tuned or clipped per FL aggregation strategy, protecting data while preserving model quality and ensuring secure FL participation.
@@
Implementation Details. The framework was implemented using Lightning [11], Flower [4], and ResNet-18 [14], and tested on an NVIDIA Tesla T4 GPU (16GB), demonstrating its feasibility in resource-constrained clinical settings. Compliance scores for each client were pre-assigned using a customizable web-based compliance scoring tool, simulating the role of a Principal Investigator (PI)(Table 2). This tool, grounded in established healthcare and security standards, evaluated clients on 12 compliance factors with predefined options and weights (Equation 1). These scores determined the level of noise dynamically added to client contributions2, ensuring baseline privacy with a minimum noise threshold applied across all clients. FL training began with the global model distributed to clients, who performed 3 epochs of training without DP on local datasets. The client contributions were then sent to the server, where noise proportional to compliance scores was added to each contribution. Before global aggregation, the server trained for one epoch on the noise-adjusted data using the aggregator dataset with DP [9]. The final aggregated model weights were computed using the selected FL strategy and redistributed to all clients. This iterative process was repeated for 50 FL training rounds, ensuring adaptive DP noise, robust aggregation, and inclusivity across clients with varying compliance levels. DP was integrated using Opacus [34], with minimum noise level tested ($101e-$ 10). Noise distribution followed the compliance score distribution, where high-compliance clients received minimal noise to preserve model performance, while low-compliance clients had higher noise applied to maintain privacy.
Table 3: Results for all combinations of Compliant Clients, Strategies, and Minimum DP Noise. Batch size is fixed at 32, and FL rounds are set to 50. Irrespective of compliance, a baseline noise of $101e-$ 10 is added to each model. Results for vanilla FL (no compliance, no DP) are included as a separate block. Detailed Experiment configurations are provided in Table 1.

### v3

@@ Abstract
Federated Learning (FL) offers a promising approach for training clinical AI models without centralizing sensitive patient data, yet its real‐world adoption is hindered by challenges in privacy, resource constraints, and compliance. Existing differential privacy (DP) approaches often apply uniform noise, which disproportionately degrades model performance even among well-compliant institutions. In this work, we propose a novel compliance-aware FL framework that enhances DP by adaptively adjusting noise based on quantifiable client compliance scores. Additionally, we introduce a compliance scoring tool based on key healthcare and security standards to promote secure, inclusive, and equitable participation across diverse clinical settings. Extensive experiments on the public datasets demonstrate that integrating under-resourced, less compliant clinics with highly regulated institutions yields accuracy improvements of up to 15% over traditional FL. This work advances FL by balancing privacy, compliance, and performance, making it a viable solution for real-world clinical workflows in global healthcare.
@@ 2 Methods
where $n$ is the total number of compliance factors, $w_{i}$ is the weight assigned to factor $i$, and $s_{i}$ is the selected option score for factor $i$. For instance, the anonymization practices factor offers three options: ISO/TS 25237:2017 Fully Anonymized (Score 1.0), Pseudonymized (Partial Anonymization) (Score 0.7), and No Anonymization (Score 0.5), with the tool defaulting to a $0.5$ threshold, adjustable by experiment owners, including setting it to 0 if needed.
@@
1:Initialize $\mathit{GLOBAL\_MODEL}$
2:for round =1=1 to $\mathit{FED\_ROUNDS}$ do
@@
5:    $\mathit{CLIENT}_{i}\leftarrow\textsc{Copy}(\mathit{GLOBAL\_MODEL})$
6:    $\mathit{CLIENT}_{i}\leftarrow\textsc{Train}(\mathit{CLIENT}_{i},\mathit{data}_{i},\mathit{epochs}=1)$
@@
11:    $\mathit{DP}_{i}\leftarrow\textsc{Copy}(\mathit{CLIENT}_{i})$
12:    $\mathit{DP}_{i}\leftarrow\textsc{DPTrain}(\mathit{DP}_{i},\mathit{agg\_data},\eta=\textsc{AdaptiveNoise}(c_{i}))$
@@
15:  $\mathit{GLOBAL\_MODEL}\leftarrow\textsc{FedAvg}(\{\mathit{DP}_{i}\})$ $\triangleright$ Fed Median/Prox/Yogi/Adam
16:  Broadcast $\mathit{GLOBAL\_MODEL}$ to clients
@@
Noise Multiplier Calculation. To implement DP adaptively, noise levels are dynamically adjusted based on client compliance scores. The noise multiplier ($N_{m}$) is computed as: $N_{m}=(1.0-S_{c})+\text{Min Noise Multiplier}$, where $S_{c}$ denotes the client’s compliance score, and Min Noise Multiplier (set to $1e\text{-}10$ in this experiment) ensures baseline privacy. This approach ensures that clients with lower compliance scores need higher noise levels. Noise can be tuned or clipped per FL aggregation strategy, protecting data while preserving model quality and ensuring secure FL participation.
@@
Implementation Details. The framework was implemented using Lightning [11], Flower [4], and ResNet-18 [14], and tested on an NVIDIA Tesla T4 GPU (16GB), demonstrating its feasibility in resource-constrained clinical settings. Compliance scores for each client were pre-assigned using a customizable web-based compliance scoring tool, simulating the role of a Principal Investigator (PI)(Table 2). This tool, grounded in established healthcare and security standards, evaluated clients on 12 compliance factors with predefined options and weights (Equation 1). These scores determined the level of noise dynamically added to client contributions2, ensuring baseline privacy with a minimum noise threshold applied across all clients. FL training began with the global model distributed to clients, who performed 3 epochs of training without DP on local datasets. The client contributions were then sent to the server, where noise proportional to compliance scores was added to each contribution. Before global aggregation, the server trained for one epoch on the noise-adjusted data using the aggregator dataset with DP [9]. The final aggregated model weights were computed using the selected FL strategy and redistributed to all clients. This iterative process was repeated for 50 FL training rounds, ensuring adaptive DP noise, robust aggregation, and inclusivity across clients with varying compliance levels. DP was integrated using Opacus [34], with minimum noise level tested (1e101e-10). Noise distribution followed the compliance score distribution, where high-compliance clients received minimal noise to preserve model performance, while low-compliance clients had higher noise applied to maintain privacy.
Table 3: Results for all combinations of Compliant Clients, Strategies, and Minimum DP Noise. Batch size is fixed at 32, and FL rounds are set to 50. Irrespective of compliance, a baseline noise of 1e101e-10 is added to each model. Results for vanilla FL (no compliance, no DP) are included as a separate block. Detailed Experiment configurations are provided in Table 1.
@@ 5 Limitations and Future Works
6 Acknowledgments
This work was supported by Grant Number 01ZZ2316A-O and partially supported by the National Cancer Institute of the National Institutes of Health under award number 5U24CA279629-02. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.